OpenStudy (anonymous):

7m+n=34 3-6n=11 substitution method

jimthompson5910 (jim_thompson5910):

7m+n=34 n=34-7m ------------ 3m-6n=11 3m-6(34-7m)=11 3m-204+42m=11 45m-204=11 45m=11+204 45m=215 m=215/45 m=43/9 n=34-7m n=34-7(43/9) n=34-301/9 n=306/9-301/9 n=5/9